OpenStudy (anonymous):

square root of (u) - square root of (v)/square root of (u) + square root of (v) =

OpenStudy (anonymous):

\[(\sqrt{u}-\sqrt{v})/(\sqrt{u}+\sqrt{v})\] ?

OpenStudy (anonymous):

Yes, that's the correctly written.

OpenStudy (anonymous):

Multiply numerator and denominator by (√u − √v ) to get rid of radicals in the denominator